A body can move along the x axis. At t = 0 it is at rest, then starts moving with acceleration a = 1.2 m/s2. What is its position after 10 seconds?
A: 120 m
B: 60 m
C: 6 m
D: 12 m
E: There are not enough data to answer
1
Expert's answer
2018-05-31T07:17:08-0400
Position of a body moving along x axis can be defined as x=x_0+v_0 t+(at^2)/2. But as we know, this body starts from the point of origin and at t=0 it is at rest, that’s why x_0=0 and v_0=0. So, equation now looks like x=(at^2)/2. After 10 seconds its position will be x=(at^2)/2=(1.2×〖10〗^2)/2=1.2×50=60 m.